The composition of a hydrated tetramethylammonium-salt of a chromium(III)-oxalate complex ion conforms to the general formula:

(N(CH3)4)xCr(C2O4)3·wH2O

A 0.4986-g sample of this compound required 36.96 mL of 0.02909 M KMnO4 solution for the titration of all of the oxalate.

Molar mass = 5.565×102 g/mol

X = 3

What is w?
Answer #1

Ans:

When ‘X=3’ the formula takes the form (N(CH3)4)3Cr(C2O4)3·wH2O

Molar mass of the complex ion, M = 5.565×102 g/mol = 556.5 g/mol

Molar mass of H2O = 18 g/mol

Molar mass of the complex ion = Molar mass of the (N(CH3)4)3Cr(C2O4)3 + (w x 18)

Molar mass of the (N(CH3)4)3Cr(C2O4)3 = 538.49 g/mol (calculate from periodic table)

Molar mass of the complex ion = Molar mass of the (N(CH3)4)3Cr(C2O4)3 + (w x 18)

556.5 = 538.49 + (w x 18)

w = 1.00055556 ~ 1

So, w = 1